OpenStudy (anonymous):

∫dx/x2+x+1 and ∫dx/(x^2 -x) are my questions. Thanks for helping.

OpenStudy (abb0t):

integrate using partial fraction decomposition

OpenStudy (anonymous):

\[1/ (x^2 + x + 1) = (Ax + B)/(x^2 + x + 1) \ \ and \ \ 1/(x^2- x) = (Ax + B)/(x^2 - x) ?\]

OpenStudy (abb0t):

Yes!

OpenStudy (abb0t):

But for the second one, you can factor out an x: \[\frac{ A }{ x } + \frac{ B }{ (x+1) }\]

OpenStudy (anonymous):

Thank you!

zepdrix (zepdrix):

Hmm I'm not sure that helps on the first one :\ We're just going to get \(\large A=0 \qquad B=1\) which is the equation we started with. Probably need to do some completing the square business on that one.